Home‎ > ‎Crew 147‎ > ‎

115-AstronomyReport

Astronomy Report
01/15/2015
Scott MacPhee
Renee Garifi

Sky Conditions: Clear
Wind Conditions: Calm
Observation Start Time: 18:00
Observation End Time: 21:00

Summary:
Tonight's observation (in progress): KC03C03248 target star with visual
magnitude of 13.28 and transit lasting 3 hours beginning at 17:59 (before
twilight). 15 Flat field images were taken at twilight along with 15 bias
frames, once the CCD camera had been sufficiently cooled.

We followed similar observing steps as last night: this time performing a
last alignment, finding the target star field, achieving good focus,
finding appropriate stars for autoguiding, and setting CCDops to grab the
series of images. We are currently collecting target 70 second light
exposures in order to capture target egress. Dark exposures will be
completed once a baseline level for the target has been established post
egress.

Objects Viewed:
Targeting KC03C03248: RA 03:34:35.87. Dec:  43:58:00.9

Problems Encountered:
Last night, we had planned to observe target KC04C016088. Due to the
lengthy set-up required on the first night of observing, we were not able
to make the targeted transit egress. We did learn to troubleshoot and
walkthrough the CCD camera operations for this telescope in order to work
more quickly tonight.

Once the temperature dropped at the end of the observation window last
evening, a distinct clink sound was heard while we were slewing the
telescope to the home position. We interpreted this sound as being due to
the internal metal motors grabbing as a result of the cold. Please advise
if there is an anomaly check procedure we can perform on the telescope.